[√n^4+9-√n^4-9]∞∑n=0 test the congervence


1
Expert's answer
2021-06-30T10:00:40-0400
"\displaystyle\sum_{n=2}^{\infin}(\sqrt{n^4+9}-\sqrt{n^4-9})"

"=\displaystyle\sum_{n=2}^{\infin}\dfrac{(\sqrt{n^4+9}-\sqrt{n^4-9})(\sqrt{n^4+9}+\sqrt{n^4-9})}{\sqrt{n^4+9}+\sqrt{n^4-9}}"

"=\displaystyle\sum_{n=2}^{\infin}\dfrac{n^4+9-n^4+9}{\sqrt{n^4+9}+\sqrt{n^4-9}}"

"=18\displaystyle\sum_{n=2}^{\infin}\dfrac{1}{\sqrt{n^4+9}+\sqrt{n^4-9}}"

The p-series "\displaystyle\sum_{n=2}^{\infin}\dfrac{1}{n^2}" converges since "p=2>1."



"\lim\limits_{n\to \infin}\dfrac{\dfrac{1}{\sqrt{n^4+9}+\sqrt{n^4-9}}}{\dfrac{1}{n^2}}"

"=\lim\limits_{n\to \infin}\dfrac{n^2}{\sqrt{n^4+9}+\sqrt{n^4-9}}"

"=\lim\limits_{n\to \infin}\dfrac{\dfrac{n^2}{n^2}}{\dfrac{\sqrt{n^4+9}+\sqrt{n^4-9}}{n^2}}"

"=\dfrac{1}{1+1}=\dfrac{1}{2}<\infin"


Therefore the given series "\displaystyle\sum_{n=2}^{\infin}(\sqrt{n^4+9}-\sqrt{n^4-9})" converges by the Limit Comparison Test.